OpenStudy (anonymous):

Solve. 2x+y=1 -3x+y=-9

OpenStudy (mertsj):

Subtract the second equation from the first equation. That will eliminate the y

OpenStudy (anonymous):

That doesn't help me much

OpenStudy (anonymous):

you can solve for y in the first equation so get y = 1- 2x. then you can plug this into the second equation to replace y so then you can solve it like -3x+(1-2x)=-9 and solve from there xD then plug the answer you get for x back into 1 - 2x to get y value

OpenStudy (anonymous):

No I know how to subtract but what do I do with what I get after subtracting?

jimthompson5910 (jim_thompson5910):

subtract equation 2 from equation 1 2x+y=1 -3x+y=-9 ---------- 5x + 0y = 10 5x = 10

jimthompson5910 (jim_thompson5910):

solve 5x = 10 for x then use this to find y

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

np